OpenStudy (anonymous):

Determine whether the lines through the pairs of points are perpendicular. A: (2,0) B: (1, -2) and C: (4,2) D: (-8, 4) I'm ok with the graphs, but I don't know which formula to use. I know that I should use the slope formula to find the slope.

OpenStudy (anonymous):

slope=\[slope(m)=(y2-y1)/(x2-x1)\]

OpenStudy (anonymous):

if m1*m2=-1 then perpendicular

OpenStudy (anonymous):

I thought it was m1= -1/m2

OpenStudy (anonymous):

its the same

OpenStudy (anonymous):

I got two different answers using both. For the one you told me about I got -1/3 = -1 and for the other formula I got 2 = 6

OpenStudy (anonymous):

slope(AB)=2 and slope(CD)=-(1/6)

OpenStudy (anonymous):

they are not perpendicular

OpenStudy (anonymous):

Thank you. @Gyan_09

OpenStudy (anonymous):

you are welcome!